If the two acute angles of a right triangle differ by 12, then what are measure of the three angles of this triangle?

Answer by mananth(16949)

let one acute agle be x
the other will be 90-x
two acute angles of a right triangle differ by 12
x-(90-x)=12
x-90+x =12
2x-90=12
2x = 102
x = 102/2 = 51
51 and 39 degrees
The two acute angles are 39 degrees and 51 degrees, and the right angle is 90 degrees.